# Dual LoRA: Enhancing LoRA with Magnitude and Direction Updates

### v1

@@ Abstract
Low-rank adaptation (LoRA) is one of the most popular methods among parameter-efficient fine-tuning (PEFT) methods to adapt pre-trained large language models (LLMs) to specific downstream tasks. However, the model trained based on LoRA often has an unsatisfactory performance due to its low-rank assumption. In this paper, we propose a novel method called Dual LoRA to improve the performance by incorporating an inductive bias into the original LoRA. Specifically, we separate low-rank matrices into two groups: the magnitude group to control whether or not and how far we should update a parameter and the direction group to decide whether this parameter should move forward or backward, to better simulate the parameter updating process of the full fine-tuning based on gradient-based optimization algorithms. We show that this can be simply achieved by adding a ReLU function to the magnitude group and a sign function to the direction group. We conduct several experiments over a wide range of NLP tasks, including natural language generation (NLG), understanding (NLU), and commonsense reasoning datasets on GPT-2, RoBERTa, DeBERTa, and LLaMA-1/2/3 as baseline models. The results show that we consistently outperform LoRA and its state-of-the-art variants with the same number of trainable parameters.
@@ 1 Introduction
Large language models (LLMs) have shown promising results on almost all natural language processing (NLP) tasks Touvron et al. (2023a); Achiam et al. (2023) and other multi-modal tasks Liu et al. (2024a), by adapting a well trained LLM to different downstream applications. Full fine-tuning (FFT) is a straightforward way to achieve this goal, but it requires tremendous computational resources and time to complete the fine-tuning process. Thus, parameter-efficient fine-tuning (PEFT) which updates a small fraction (less than $2\%$) of parameters has attracted more and more attention due to its low memory and time requirements.
@@
Traditional PEFT methods include adapter tuning Hu et al. (2023) which adds trainable tiny modules to adapt to downstream tasks, prompt tuning Peng et al. (2024) that inserts learnable prompt vectors to the existing input, and low-rank adaptation (LoRA) Hu et al. (2021a) which updates the original parameters by adding low-rank matrices. Among them, LoRA surpasses other methods by achieving better performance without generating additional inference costs.
@@
Many follow-ups manage to improve the fine-tuning performance of LoRA. LoRA+ Hayou et al. (2024) uses different learning rates to update low-rank matrices $A$ and $B$ and enhance the performance with a well-chosen learning rate ratio. DoRA Liu et al. (2024b) decomposes the original weight matrix into a normalized matrix and its corresponding norm and applies the original LoRA to the normalized matrix. FLoRA Si et al. (2024) generates LoRA to high dimensional space and inserts a low-rank core matrix into the original LoRA matrices to improve its performance. MoRA Jiang et al. (2024) replaces the low-rank matrices with a square matrix to achieve high-rank updating and applies a compress layer and a decompress layer to maintain a roughly similar number of trainable parameters. However, they share a common drawback: as the trainable parameters are much fewer than those of FFT, updating them without incorporating prior knowledge will inevitably result in unsatisfactory model accuracy.
@@
Thus, in this paper we introduce an inductive bias into the original LoRA method, i.e., to simulate the parameter updating process of FFT, which utilizes gradient-based optimization algorithms. Specifically, we divide the low-rank matrices into two groups: the magnitude group, which controls whether and to what extent a parameter should be updated; and the direction group, which determines the direction of the update—–whether it should be positive or negative. The whole fine-tuning process can be treated as adjusting the sign and magnitude of each element in the update matrix and adding them back to the original parameters to gradually achieve the optimal solution. We conduct experiments to validate the effectiveness of our method over a wide range of NLP tasks including natural language generation (NLG), understanding (NLU), and commonsense reasoning to make a fair comparison with state-of-the-art methods. The evaluation results on different LLM models such as GPT-2, RoBERTa, DeBERTa, LLaMA-7B/13B, LLaMA2-7B, LLaMA3-8B, and LLaMA3-70B-Instruct show that we can achieve consistent improvements over these SOTA methods by using the same number of training parameters.
@@ 2.1 PEFT Methods in LLMs
Prefix tuning is the first kind of methods Li and Liang (2021); Liu et al. (2022); Zhang et al. (2024) in PEFT. It was first proposed by Li et.al. Li and Liang (2021), which was a lightweight alternative to FFT that kept LLM parameters frozen and only optimized a sequence of continuous task-specific vectors called prefix. Dynamic prefix-tuning Liu et al. (2022) proposed a generative template-based event extraction method with dynamic prefixes by integrating context information with type-specific prefixes to learn a context-specific prefix for each context. Selective prefix-tuning Zhang et al. (2024) showed that prefix tokens carried context-specific information and enhanced their specialization can improve model performance. Thus, they integrated a selective mechanism inspired by selective self-attention and introduced selective loss to encourage diversity in prefix tokens.
Prompt tuning is the second kind of PEFT method that added trainable embeddings to original word embeddings and learned these soft prompts through back-propagation and tuned them to incorporate signals from any number of labeled examples Lester et al. (2021). P-Tuning v2 Liu et al. (2021) empirically found that properly optimized prompt tuning can be universally effective across a wide range of model scales and NLU tasks, which increased the capacity of continuous prompts and closed the gap to FFT. Knowledgeable Prompt-tuning Hu et al. (2021b) improved and stabilized the original prompt-tuning method by expanding the label word space of the verbalizer with external knowledge bases and refining it with PLM before predicting.
Representation fine-tuning (REFT) aims to train interventions that manipulate model representations to steer model behaviors on downstream tasks at inference time. ReFT Wu et al. (2024) introduced a family of ReFT methods that operated on a frozen base model and learned task-specific interventions on hidden representations.
@@ 2.2 LoRA-Based Methods
LoRA Hu et al. (2021a) assumed that only a small number of task-specific parameters needed to be tuned to fit the downstream tasks and updated the weights with two low-rank matrices. These matrices can be merged back into the original weights during inference to avoid additional computational costs. LoRA+ Hayou et al. (2024) argued that LoRA led to sub-optimal results, and the problem can be corrected by setting different learning rates for the low-rank matrices $A$ and $B$ with a fixed learning rate ratio. MoRA Jiang et al. (2024) believed that the low-rank updating mechanism limited the ability of LLMs and used a square matrix to achieve high-rank updating with the same number of trainable parameters. Two non-parameter operators were used to reduce the input dimension and increase the output dimension of this square matrix. DoRA Liu et al. (2024b) decomposed the pre-trained weight into magnitude and direction for fine-tuning, and employed original LoRA for direction component update to accelerate the training process.
The methods mentioned above can improve the performance of downstream tasks. However, their performance is still unsatisfactory because of the low-rank assumption Hu et al. (2021a); Hayou et al. (2024); Liu et al. (2024b). Although MoRA Jiang et al. (2024) attempted to address this issue by using a high-rank matrix, its rank and the number of trainable parameters remained significantly lower than those in FFT. Thus, it is difficult to achieve satisfactory model performance without incorporating prior knowledge into the training process.
@@
Another family of methods aim to modify the gradient calculation and backward propagation process of training, such as GaLore Zhao et al. (2024), FLoRA Hao et al. (2024) and GaRare Liu et al. . These methods are orthogonal to the proposed Dual LoRA which only focuses on the architecture and forward pass modification, and a detailed discussion falls outside the scope of this paper.
@@ 3.1 Low-Rank Adaptation (LoRA)
Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA Hu et al. (2021a) assumes that a low “intrinsic rank" is enough during adaptation on downstream tasks and constrains the updated matrix with a low-rank decomposition:
@@ 3.2 Optimization Methods
To achieve this, different optimization methods leverage different ways to compute $\Delta w$. For example, gradient descent Bottou (2010) uses $\Delta w=\frac{1}{n}\sum_{i=1}^{n}\nabla_{w}\ell(f(x_{i}),y_{i})$ to compute the update, and Adam Kingma (2014) uses $\Delta w=\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)$ where $\hat{m}_{t}$ and $\hat{v}_{t}$ are the first-moment estimate and second-moment estimate, and $\epsilon=10^{-8}.$
@@ 3.3 Dual LoRA
where $\rm Sign(\cdot)$} is an element-wise operation that outputs +1+1 for positive input and 1-1 otherwise. Note that the gradient of the sign function is zero almost everywhere, and backward propagation cannot be applied during training. Thus, given $x_{b}=Sign(x)$, the straight-through estimator (STE) method Bengio et al. (2013) is introduced to compute its gradient as:
@@
LoRA ($r=64$) 1.64 67.3 79.0 76.3 76.6 78.8 74.5 59.3 77.4 73.6
DoRA ($r=32$) 0.84 68.7 83.3 79.4 85.5 81.3 80.8 66.0 78.8 78.0
DoRA ($r=64$) 1.65 68.9 82.1 77.4 75.9 80.0 80.0 64.8 81.0 76.3
Dual LoRA ($r=32$) 1.64 70.0 83.2 79.5 87.3 83.0 81.6 65.2 81.0 78.9
@@
DoRA ($r=16$) 0.43 63.5 82.8 79.5 90.6 82.4 83.9 69.9 81.8 79.3
DoRA ($r=32$) 0.84 72.2 83.5 80.3 89.0 82.5 84.1 69.5 80.4 80.2
Dual LoRA ($r=16$) 0.83 72.3 83.3 79.8 89.8 84.6 84.8 70.2 82.8 81.0
@@
DoRA ($r=16$) 0.35 75.1 87.8 80.8 95.6 86.3 90.4 80.0 85.6 85.2
DoRA ($r=32$) 0.71 71.7 88.0 80.2 95.5 86.6 90.7 78.4 85.0 84.5
Dual LoRA ($r=16$) 0.70 75.5 89.2 81.4 95.8 86.0 90.5 81.1 86.6 85.8
@@ 4 Experiments
In this section, we evaluate the effectiveness of the proposed Dual LoRA on various NLP tasks. We compare our methods with other PEFT competitors by fine-tuning LlaMA-7B/13B, LLaMA2-7B, LLaMA3-8B, and LLaMA3-70B-Instruct models on a series of commonsense reasoning datasets. Then, we explore the ability of our method on the neural language understanding (NLU) dataset GLUE by fine-tuning RoBERTa base/large and DeBERTa XXL. Furthermore, we conduct experiments on neural language generation (NLG) datasets including E2E NLG Challenge, DART and WebNLG using GPT2_M and GPT2_L as backbones (see Appendix A). All experiments above show that Dual LoRA can surpass other LoRA-based methods with the same or fewer trainable parameters and achieve state-of-the-art results. Finally, we analyze our method further by performing a series of ablation studies. In the following experiments, we set the rank of the magnitude group and the direction group as the same, i.e., $r_{1}=r_{2}=r$ unless specified.
Competitors. We compare Dual LoRA with a series of baseline methods including LoRA-based methods (LoRA Hu et al. (2021a), LoRA+ Hayou et al. (2024), GaLore Zhao et al. (2024), GaRare Liu et al. , Delta-LoRA Zi et al. (2023), CorDA Yang et al. (2024), VeRA Kopiczko et al. (2024), RandLoRA Albert et al. (2025), and DoRA Liu et al. (2024b)) and other PEFT methods (efficient adapter design with LayerNorm (Adapter-L) Lin et al. (2020), parallel adapter tuning (Adapter-P) He et al. (2021) and prefix-layer tuning (Prefix) Li and Liang (2021)).
@@ 4.1 Commonsense Reasoning
Datasets and baseline models. We evaluate Dual LoRA and different PEFT methods on the commonsense reasoning task which is composed of eight different sub-tasks including BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), Social IQa Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy/challange Clark et al. (2018) and OpenBookQA Mihaylov et al. (2018). Similarly to DoRA, we merge the training sets from all sub-tasks to get the final training set and perform evaluations on their own testing datasets for each task.
For the baseline models, we use LLaMA-7B/13B Touvron et al. (2023a), LLaMA2-7B Touvron et al. (2023b), LLaMA3-8B Dubey et al. (2024), and LLaMA3-70B-Instruct Dubey et al. (2024). We halve the rank of our low-rank matrices to ensure that the same number of trainable parameters are used compared to other LoRA-based methods. We tune the learning rate for our method, and all other training hyper-parameters are kept unchanged as in DoRA in order to make a fair comparison. We train one epoch for LLaMA3-70B-instruct, and three epochs for other baseline models.
Results. The results in Tab. 1 show that we can consistently outperform DoRA with less trainable parameters on all of the baseline models. For example, Dual LoRA enhances the average accuracy by 0.9%/0.6% compared to the previous best result on LLaMA-7B/13B. The performance gains are still notable on LLaMA2-7B, LLaMA3-8B and LLaMA3-70B-Instruct, which are 0.8%, 0.2% and 0.8%.

### v3

@@ Abstract
Low-rank adaptation (LoRA) is one of the most popular methods among parameter-efficient fine-tuning (PEFT) methods to adapt pre-trained large language models (LLMs) to specific downstream tasks. However, the model trained based on LoRA often has an unsatisfactory performance due to its low-rank assumption. In this paper, we propose a novel method called Dual LoRA to improve the performance by incorporating an inductive bias into the original LoRA. Specifically, we separate low-rank matrices into two groups: the magnitude group to control whether or not and how far we should update a parameter and the direction group to decide whether this parameter should move forward or backward, to better simulate the parameter updating process of the full fine-tuning based on gradient-based optimization algorithms. We show that this can be simply achieved by adding a ReLU function to the magnitude group and a sign function to the direction group. We conduct several experiments over a wide range of NLP tasks, including natural language understanding (NLU) and commonsense reasoning datasets on RoBERTa, DeBERTa, and LLaMA-1/2/3 as baseline models. The results show that we consistently outperform LoRA and its state-of-the-art variants with the same number of trainable parameters.
@@ 1 Introduction
Large language models (LLMs) have shown promising results on almost all natural language processing (NLP) tasks (Touvron et al., 2023a; Achiam et al., 2023) and other multi-modal tasks (Liu et al., 2024a), by adapting a well trained LLM to different downstream applications. Full fine-tuning (FFT) is a straightforward way to achieve this goal, but it requires tremendous computational resources and time to complete the fine-tuning process. Thus, parameter-efficient fine-tuning (PEFT) which updates a small fraction (less than $2\%$) of parameters has attracted more and more attention due to its low memory and time requirements.
@@
Traditional PEFT methods include adapter tuning (Hu et al., 2023) which adds trainable tiny modules to adapt to downstream tasks, prompt tuning (Peng et al., 2024) that inserts learnable prompt vectors to the existing input, and low-rank adaptation (LoRA) (Hu et al., 2021a) which updates the original parameters by adding low-rank matrices. Among them, LoRA surpasses other methods by achieving better performance without generating additional inference costs.
@@
Many follow-ups manage to improve the fine-tuning performance of LoRA. LoRA+ (Hayou et al., 2024) uses different learning rates to update low-rank matrices $A$ and $B$ and enhance the performance with a well-chosen learning rate ratio. DoRA (Liu et al., 2024b) decomposes the original weight matrix into a normalized matrix and its corresponding norm and applies the original LoRA to the normalized matrix. FLoRA (Si et al., 2024) generates LoRA to high dimensional space and inserts a low-rank core matrix into the original LoRA matrices to improve its performance. MoRA (Jiang et al., 2024) replaces the low-rank matrices with a square matrix to achieve high-rank updating and applies a compress layer and a decompress layer to maintain a roughly similar number of trainable parameters. However, they share a common drawback: as the trainable parameters are much fewer than those of FFT, updating them without incorporating prior knowledge will inevitably result in unsatisfactory model accuracy.
@@
Thus, in this paper we introduce an inductive bias into the original LoRA method, i.e., to simulate the parameter updating process of FFT, which utilizes gradient-based optimization algorithms. Specifically, we divide the low-rank matrices into two groups: the magnitude group, which controls whether and to what extent a parameter should be updated; and the direction group, which determines the direction of the update—–whether it should be positive or negative. The whole fine-tuning process can be treated as adjusting the sign and magnitude of each element in the update matrix and adding them back to the original parameters to gradually achieve the optimal solution. We conduct experiments to validate the effectiveness of our method over a wide range of NLP tasks including natural language understanding (NLU) and commonsense reasoning to make a fair comparison with state-of-the-art methods. The evaluation results on different LLM models such as RoBERTa, DeBERTa, LLaMA-7B/13B, LLaMA2-7B, LLaMA3-8B, and LLaMA3-70B-Instruct show that we can achieve consistent improvements over these SOTA methods by using the same number of training parameters.
@@ 2.1 PEFT Methods in LLMs
Prefix tuning is the first kind of methods (Li and Liang, 2021; Liu et al., 2022; Zhang et al., 2024) in PEFT. It was first proposed by Li et.al. (Li and Liang, 2021), which was a lightweight alternative to FFT that kept LLM parameters frozen and only optimized a sequence of continuous task-specific vectors called prefix. Dynamic prefix-tuning (Liu et al., 2022) proposed a generative template-based event extraction method with dynamic prefixes by integrating context information with type-specific prefixes to learn a context-specific prefix for each context. Selective prefix-tuning (Zhang et al., 2024) showed that prefix tokens carried context-specific information and enhanced their specialization can improve model performance. Thus, they integrated a selective mechanism inspired by selective self-attention and introduced selective loss to encourage diversity in prefix tokens.
Prompt tuning is the second kind of PEFT method that added trainable embeddings to original word embeddings and learned these soft prompts through back-propagation and tuned them to incorporate signals from any number of labeled examples (Lester et al., 2021). P-Tuning v2 (Liu et al., 2021) empirically found that properly optimized prompt tuning can be universally effective across a wide range of model scales and NLU tasks, which increased the capacity of continuous prompts and closed the gap to FFT. Knowledgeable Prompt-tuning (Hu et al., 2021b) improved and stabilized the original prompt-tuning method by expanding the label word space of the verbalizer with external knowledge bases and refining it with PLM before predicting.
Representation fine-tuning (REFT) aims to train interventions that manipulate model representations to steer model behaviors on downstream tasks at inference time. ReFT (Wu et al., 2024) introduced a family of ReFT methods that operated on a frozen base model and learned task-specific interventions on hidden representations.
@@ 2.2 LoRA-Based Methods
LoRA (Hu et al., 2021a) assumed that only a small number of task-specific parameters needed to be tuned to fit the downstream tasks and updated the weights with two low-rank matrices. These matrices can be merged back into the original weights during inference to avoid additional computational costs. LoRA+ (Hayou et al., 2024) argued that LoRA led to sub-optimal results, and the problem can be corrected by setting different learning rates for the low-rank matrices $A$ and $B$ with a fixed learning rate ratio. MoRA (Jiang et al., 2024) believed that the low-rank updating mechanism limited the ability of LLMs and used a square matrix to achieve high-rank updating with the same number of trainable parameters. Two non-parameter operators were used to reduce the input dimension and increase the output dimension of this square matrix. DoRA (Liu et al., 2024b) decomposed the pre-trained weight into magnitude and direction for fine-tuning, and employed original LoRA for direction component update to accelerate the training process.
The methods mentioned above can improve the performance of downstream tasks. However, their performance is still unsatisfactory because of the low-rank assumption (Hu et al., 2021a; Hayou et al., 2024; Liu et al., 2024b). Although MoRA (Jiang et al., 2024) attempted to address this issue by using a high-rank matrix, its rank and the number of trainable parameters remained significantly lower than those in FFT. Thus, it is difficult to achieve satisfactory model performance without incorporating prior knowledge into the training process.
@@
Another family of methods aim to modify the gradient calculation and backward propagation process of training, such as GaLore (Zhao et al., 2024), FLoRA (Hao et al., 2024) and GaRare (Liu et al., ). These methods are orthogonal to the proposed Dual LoRA which only focuses on the architecture and forward pass modification, and a detailed discussion falls outside the scope of this paper.
@@ 3.1 Low-Rank Adaptation (LoRA)
Given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA (Hu et al., 2021a) assumes that a low “intrinsic rank" is enough during adaptation on downstream tasks and constrains the updated matrix with a low-rank decomposition:
@@ 3.2 Optimization Methods
To achieve this, different optimization methods leverage different ways to compute $\Delta w$. For example, gradient descent (Bottou, 2010) uses $\Delta w=\frac{1}{n}\sum_{i=1}^{n}\nabla_{w}\ell(f(x_{i}),y_{i})$ to compute the update, and Adam (Kingma, 2014) uses $\Delta w=\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)$ where $\hat{m}_{t}$ and $\hat{v}_{t}$ are the first-moment estimate and second-moment estimate, and $\epsilon=10^{-8}.$
@@ 3.3 Dual LoRA
where $\rm Sign(\cdot)$} is an element-wise operation that outputs +1+1 for positive input and 1-1 otherwise. Note that the gradient of the sign function is zero almost everywhere, and backward propagation cannot be applied during training. Thus, given $x_{b}=Sign(x)$, the straight-through estimator (STE) method (Bengio et al., 2013) is introduced to compute its gradient as:
@@
LoRA ($r=32$) 0.83 67.5 80.8 78.2 83.4 80.4 78.0 62.6 79.1 76.3
DoRA ($r=32$) 0.84 69.7 83.4 78.6 87.2 81.0 81.9 66.2 79.2 78.4
DoRA ($r=64$) 1.65 70.1 82.0 75.6 85.9 79.7 79.1 63.7 78.4 76.8
Dual LoRA ($r=16$) 0.83 71.3 83.4 81.0 88.1 83.7 85.0 71.0 82.4 80.7
@@
DoRA ($r=16$) 0.43 72.0 83.1 79.9 89.1 83.0 84.5 71.0 81.2 80.5
DoRA ($r=32$) 0.84 71.8 83.7 76.0 89.1 82.6 83.7 68.2 82.4 79.7
HiRA ($r=32$) 0.83 71.2 83.4 79.5 88.1 84.0 86.7 73.8 84.6 81.4
Dual LoRA ($r=16$) 0.83 73.8 83.8 80.3 88.1 85.2 87.7 74.1 85.0 82.3
@@
DoRA ($r=16$) 0.35 74.5 88.8 80.3 95.5 84.7 90.1 79.1 87.2 85.0
DoRA ($r=32$) 0.71 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
HiRA ($r=32$) 0.70 75.4 89.7 81.2 95.4 87.7 93.3 82.9 88.3 86.7
Dual LoRA ($r=16$) 0.70 76.2 89.1 83.2 96.0 89.1 93.7 85.4 87.6 87.5
@@ 4 Experiments
In this section, we evaluate the effectiveness of the proposed Dual LoRA on various NLP tasks. We compare our methods with other PEFT competitors by fine-tuning LlaMA-7B/13B, LLaMA2-7B, LLaMA3-8B, and LLaMA3-70B-Instruct models on a series of commonsense reasoning datasets. Then, we explore the ability of our method on the neural language understanding (NLU) dataset GLUE by fine-tuning RoBERTa base/large and DeBERTa XXL. Furthermore, we conduct experiments on neural language generation (NLG) with E2E NLG Challenge dataset using GPT2_M and GPT2_L as backbones (see Appendix A). All experiments above show that Dual LoRA can surpass other LoRA-based methods with the same or fewer trainable parameters and achieve state-of-the-art results. Finally, we analyze our method further by performing a series of ablation studies. In the following experiments, we set the rank of the magnitude group and the direction group as the same, i.e., $r_{1}=r_{2}=r$ unless specified.
Competitors. We compare Dual LoRA with a series of baseline methods including LoRA-based methods (LoRA (Hu et al., 2021a), LoRA+ (Hayou et al., 2024), GaLore (Zhao et al., 2024), GaRare (Liu et al., ), Delta-LoRA (Zi et al., 2023), CorDA (Yang et al., 2024), VeRA (Kopiczko et al., 2024), RandLoRA (Albert et al., 2025), DoRA (Liu et al., 2024b), and HiRA (huang2025hira)) and other PEFT methods (efficient adapter design with LayerNorm (Adapter-L) (Lin et al., 2020), parallel adapter tuning (Adapter-P) (He et al., 2021) and prefix-layer tuning (Prefix) (Li and Liang, 2021)).
@@ 4.1 Commonsense Reasoning
Datasets and baseline models. We evaluate Dual LoRA and different PEFT methods on the commonsense reasoning task which is composed of eight different sub-tasks including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), Social IQa (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-easy/challange (Clark et al., 2018) and OpenBookQA (Mihaylov et al., 2018). Similarly to DoRA, we merge the training sets from all sub-tasks to get the final training set and perform evaluations on their own testing datasets for each task.
For the baseline models, we use LLaMA-7B/13B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), LLaMA3-8B (Dubey et al., 2024), and LLaMA3-70B-Instruct (Dubey et al., 2024). We halve the rank of our low-rank matrices to ensure that the same number of trainable parameters are used compared to other LoRA-based methods. We tune the learning rate for our method, and all other training hyper-parameters are kept unchanged as in HiRA in order to make a fair comparison. We train one epoch for LLaMA3-70B-instruct, and three epochs for other baseline models.
Results. The results in Tab. 1 show that we can consistently outperform DoRA and HiRA with less trainable parameters on all of the baseline models. For example, Dual LoRA enhances the average accuracy by 2.3%/3.1% compared to the previous best result on LLaMA-7B/13B. The performance gains are still notable on LLaMA2-7B, LLaMA3-8B and LLaMA3-70B-Instruct, which are 0.9%, 0.8% and 1.9%.